OpenStudy (anonymous):

I don't know where to start with this problem. Not asking for it to be solved. sin(theta + pie/4)sec 2theta in terms of theta alone

OpenStudy (anonymous):

woah

OpenStudy (mathmale):

Your sin(theta + Pi/4) immediately brings to mind the sum formula \[\sin(a+b)=\sin a \cos b + \cos a \sin b.\] Try that, using a =theta and b = Pi/4.

OpenStudy (mathmale):

I'll be back later.

OpenStudy (anonymous):

ok thank u, will try that...

OpenStudy (anonymous):

Thanks math male, that was really helpful. \[\sqrt{2}/2(\cos \theta - \sin \theta)\]